# GLEE: A Unified Framework and Benchmark for Language-based Economic Environments

### v1

@@ 2.3 Persuasion games
We consider the following degrees of freedom in persuasion games: (i) whether or not Alice knows Bob’s high-quality product valuation $v$ (complete vs. incomplete information); (ii) whether or not the players know when the game terminates (finite vs. infinite); (iii) whether or not Alice’s messages involve natural language (structured vs. linguistic); (iv) the values of $v$, $p$ and $M$; (v) the value of $T$ in the finite horizon case; (vi) whether the game is with a long-living Bob or with myopic buyers.
@@
3 Data collection
In this section, we describe the data collection process. We developed a game management system to facilitate data collection from the games described in §2. The system is written in Python and is easy to use and customize, allowing future researchers to collect data seamlessly. New language models can be added easily, enabling them to play in any configurable setup. With a single command line, the system allows running either a single configuration or a set of configurations. Prompt samples are described in Appendix B.
3.1 LLM data collection
@@
In total 384 configurations In total 576 configurations In total 384 configurations
Table 1: Parameters and their optional values used to define the 1,344 game configurations across bargaining, negotiation, and persuasion game families for data collection. $T=\infty$ means a very large value of $T$, which is unknown to the players. CI = Complete Information. MA = Textual messages allowed.
human Qwen-2 Gem Lam-3 Lam-3.1 Barg. Nego. Pers. total
# Games 3.4K 309K 429K 342K 332K 282K 539K 133K 954K
# Messages 6.64K 1.16M 870K 1.34M 1.33M  [^6] K 1.41M 4.01M 6.09M
# Words 58.8K 40.9M 29.2M 51.7M 54.7M 10.2M 52M 174M 236M
# Decision 7.43K 1.42M 1.11M 1.58M 1.56M 1.01M 2.15M 4M 7.15M
Table 2: Statistics of data collected: by human players and each of the LLMs (left), for each game family (center), and in total (right). The number of games indicates the number of games in which the language model was involved as one or both players. The number of messages and decisions represents the number of decisions made and messages sent by the language model. Gem. = Gemini, Lam = Llama; Barg. = Bargening, Nego. = Negotiation, Pers. = Persuasion.
@@
Since the game space defined by the parameters presented in §2 is infinite for each of the game families, it is clear that data cannot be collected from all possible games. Therefore, we attempted to cover the game space by selecting diverse values for each of the parameters defining the games, and we collected data from every possible combination generated by these parameters.
Table 1 shows the parameters defining the groups from which we collected data. In total, we collected data from 1,344 different configurations: 384 configurations of bargaining, 576 configurations of negotiation, and 384 configurations of persuasion games.
For data collection, we used four language models: Qwen-2-7B (Yang et al. 2024), Llama-3-8B, Llama-3.1-8B (Meta 2024), and Gemini-1.5-Flash (Gemini Team 2024). Each of the 16 possible pairs (including a language model against itself, with attention to the order defining the roles of the players in the game) played at least 30 games from the 1,344 configurations. In total, 954K games were played, with full statistics appearing in Table 2. The complete data from all games, including the messages sent by the language models, the decisions made, and the prompts and responses of the language models that generated these messages, form one of the main contributions of this paper and are available in the paper’s GitHub repository.
In addition to these models, we collected data for some of the configurations using Gemini-1.5-Pro (Gemini Team 2024) and Llama-3.1-405B (Meta 2024), two larger, more powerful, and expensive models compared to the ones used. These models played against each other and against some of the four presented models. Due to budget constraints, data collection with these models was not comprehensive, but the data from these games still contributes to the paper.
3.2 Human Data Collection
One of the main objectives of collecting data from language games is to compare the behavior of LLMs with human behavior in economic and strategic situations. To facilitate this comparison, we developed an interface that allows human players to play all the language games that can be defined using GLEE. The interface transforms the various prompts presented to LLMs into user-friendly screens for human participants, displaying the prompt and requesting them to send messages and make decisions. [^12] Since human players are accustomed to being addressed by their first name (rather than as Alice or Bob), we asked them to enter their name at the beginning of the game and referred to them by their name throughout the game. The player’s name was the only difference between the human player and the LLM-based player. Through this interface, we enable human players to take on the role of one of the players while the other player is a pre-selected LLM. The interface is one of the contributions of this paper. Screenshots of the interface can be found in Appendix C.1.
The interface, developed using oTree (Chen et al. 2016), enables integration with Amazon’s crowdsourcing platform, mTurk, [^13] https://www.mturk.com/. through which we recruited 3,405 players who participated in various configurations against Gemini-1.5-flash. We chose Gemini-1.5-flash since it demonstrated strong performance compared to other LLMs and allowed comprehensive data collection due to its low usage cost. Since we aimed to collect multiple games from each configuration, we had to select a limited set of configurations for human participants to play. The process of selecting these sets is described in Appendix C.2. We collected human data from 195 different configurations: 78 of which were bargaining games, 60 of which were negotiation games, and 57 were persuasion games.
Each human player was allowed to play one game every 12 hours from each family of games. Human players were paid a base rate calculated at \$6 per hour, plus an average bonus of \$6 per hour. In total, we paid \$2,245 to all players for their participation in the games. Bonuses were dependent on the configuration the human players played and their success in the game. The average bonus was known to players at the start of the game, aiming to encourage serious gameplay. To ensure players remained focused and made thoughtful decisions, we conducted two attention checks during the experiment, detailed in Appendix C.3. Players who failed in one of the attention checks were excluded from the dataset. To reflect the real-world significance of the magnitude of product prices (the parameter M) in each family of games, we defined the bonus as dependent on M: in configurations where M = $10^{2}$, the average bonus was \$3 per hour; where M = $10^{4}$, the average bonus was \$6 per hour; and where M = $10^{6}$, the average bonus was \$9 per hour.
4 Exploratory Data Analysis
In this section, we first present a method to ensure an adequate comparison between models that played different configurations. Then, we address two key research questions: Q1: How did the language models perform in each game family in terms of self-gain, efficiency, and fairness? Q2: How did the humans perform compared to the language models? Q3: How do the parameters defining the games influence the various metrics?
Adequate Comparison Between Models
For the four language models introduced in §3, we collected data covering the entire set of the configurations defined in Table 1 for every possible pair of language models. This enables seamless comparison of their performance across all defined metrics. Henceforth, this dataset will be referred to as the baseline set, and the models that generated it will be referred to as the baseline models.
When evaluating data generated by a new model or by humans, it is often not feasible (due to cost considerations) to have them play all games against all existing language models. [^14] Collecting data from humans playing 30 games in each possible configuration would cost approximately \$187K. Collecting similar data from Gemini-1.5-Pro would cost around \$10K. To evaluate model performance on a large variety of language games, we can collect data from a subset of configurations, calculate the metrics for those, and use the results (along with the metrics from the baseline set) to train a predictive model that estimates the metrics for configurations for which data was not collected. This method allows for an adequate comparison between models that played different configuration sets.
For the calculation of the metrics, we train a simple regression model for each metric. The model receives as input the values defining the configuration as binary features (e.g., ”Is $M=100$?”) and binary features describing the interaction between the language models and other parameter values (e.g., ”Did Alice played by Gemini-1.5-Flash and $M=100$?” and ”Did Alice played by Gemini-1.5-Flash, Bob played by Llama-3.1-8B, and $M=100$?”). To prevent the model from learning relationships between parameter values directly from configuration settings, we used binary features even for parameters that contain numerical values. We trained the models on the data from all games and predicted the values for all configurations: both those for which data was collected and those for which it was not. Therefore, to maintain consistency, all the metrics are calculated by averaging the predicted values over all game configurations. This approach allows us to reasonably estimate the metric values expected for configurations where no human data was collected. The exact formulas defining each regression model are presented in Appendix D.
To assess the quality of the models, we examined the adjusted R-squared values of the regression models. In 10 out of 12 trained models, we obtained adjusted R-squared values higher than 0.38. Table 3 shows the adjusted R-squared values of the models. The table shows that, in general, persuasion games are the easiest to predict, and we hypothesize that this is because their outcomes consist of a collection of multiple decisions. Negotiation games are the most difficult to predict, possibly because, unlike the other games, players could exit the game at any given moment. It seems that fairness is the most challenging metric to fit with a regression model, with negotiation games showing an especially low adjusted R-squared.
Alice gain Bob gain Efficiency Fairness
Bargaining 0.54 0.47 0.57 0.28
Negotiation 0.48 0.58 0.38 0.09
Persuasion 0.63 0.89 0.58 0.41
Table 3: Adjusted R-square values of the regression models used for metric calculations and analyzing the impact of specific parameters on metrics.
Family Rule Metric Human Qwen-2 Lam-3 Lam-3.1 Gem-F
Bargaining Alice Self Gain 0.61 0.49 0.43 0.46 0.55
Efficiency 0.89 0.88 0.86 0.86 0.88
Fairness 0.71 0.87 0.87 0.86 0.76
Bob Self Gain 0.27 0.34 0.38 0.42 0.43
Efficiency 0.94 0.97 0.79 0.78 0.88
Fairness 0.69 0.81 0.86 0.80 0.91
Negotiation Alice Self Gain -0.24 -0.02 0.03 0.05 0.07
Efficiency 0.65 0.68 0.75 0.73 0.75
Fairness 0.39 0.82 0.89 0.91 0.88
Bob Self Gain -0.05 0.08 0.05 0.07 0.12
Efficiency 0.67 0.73 0.67 0.69 0.81
Fairness 0.58 0.81 0.83 0.84 0.82
Persuasion Alice Self Gain 0.55 0.60 0.57 0.61 0.58
Efficiency 0.55 0.78 0.71 0.76 0.65
Fairness 0.41 0.63 0.61 0.60 0.53
Bob Self Gain 0.68 0.45 0.53 0.75 0.41
Efficiency 0.93 0.84 0.68 0.35 0.65
Fairness 0.21 0.59 0.60 0.61 0.77
Table 4: Average predicted self-gain, efficiency, and fairness metrics for each language model across different roles and game families, comparing baseline models and human players. In bold: The highest value achieved in the metric.
Model and Human Performance (Q1 and Q2)
Table 4 presents, for each language model, the metric values for each role in each game family. The presented values are the predicted average values over all configurations for which data was collected, as described in §3. For each role, we present the self-gain achieved by the model, as well as the two metrics introduced in §2: the efficiency and fairness values observed in games involving the model. Note that while the self-gain metric measures the player’s performance in the game, the efficiency and fairness metrics describe the economic quality of the interaction outcome. This table allows us to compare the performance of different language models across game families.
The table allows us to compare the performance of humans to that of language models across different game families. Humans outperformed the language models in bargaining games when playing as Alice, but when playing as Bob, they achieved the worst performance. In negotiation games, humans performed poorly, with a negative average self gain, in contrast to the language models, which managed to achieve gains in both roles (except for Qwen-2 in the role of Alice). Human performance in persuasion games was relatively strong, second only to that of Llama-3.1.
@@ 4 Exploratory Data Analysis
Importantly, metric values presented in Table 4 are calculated by averaging over all possible game configurations in our dataset. These averages are therefore highly sensitive to the particular configurations, and the configuration distributions, in our dataset. To tailor the benchmark to specific applications, we recommend re-defining the parameter space and their distributions according to the economic context. [^15] For instance, one could ask whether agents’ performance significantly differs in economic environments where inflation is high (translating into lower discount factors, in bargaining games). To evaluate such a scenario, one can simulate games in which discount factor distribution fits these conditions, and re-evaluate agents’ performance with respect to the new distribution of configurations. In addition, notice that the fact that metrics are averaged over all configurations may hide the effect of particular game parameters on the game outcome or agents’ self-gain. We next demonstrate how a more delicate analysis of the dataset enables to isolation of the effect of certain game parameters.
@@
To evaluate the effect of a specific parameter on a metric, we use the same regression model. To create a graph depicting the effect of a parameter X on a metric Y achieved by a language model Z playing as Alice, we extract from the regression model the effects of X values, Z values, and their interaction on Y, summing the relevant values. The resulting depiction shows the impact of parameter changes relative to a baseline configuration defined by one of X’s values. Appendix E presents graphs describing the impact of each parameter on player gains, fairness, and efficiency in games, depending on the language model that played the games. The graphs include confidence intervals at a 95% confidence level.
Figure 1 serve as example from which conclusions can be drawn about the impact of the parameters. Figure 1 (Right) shows that full information reduces efficiency for all models in bargaining games (and also in negotiation games, as shown by Figure 24 in Appendix E). While this may be counter-intuitive at first, it might be the case that complete information prevents trade when efficient trade is possible. [^16] Suppose, for example, that Alice values the product at 80\$ and Bob values the product at 100\$. If Alice sets a price of 99\$, and Bob knows Alice’s valuation, then Bob might reject the offer even though it is beneficial to him, since he feels that Alice should have made a more decent offer that takes into consideration Bob’s satisfaction as well. Figure 1 (Left) shows that sending textual messages increases efficiency in negotiation games (a trend that also appears in bargaining games, as shown by Figure 19 in Appendix E). In persuasion games, sending textual messages improves Alice’s gains for all models (as can be seen in Figure 31 in Appendix E), a result observed by Apel et al. (2022) in language-based games involving humans.
Figure 1: Right: the negative effect that full information has on efficiency in the bargaining game, depending on the player model. Left: the positive effect that sending textual messages has in negotiation games, depending on the player model.

### v2

@@ 2.3 Persuasion Games
We consider the following degrees of freedom in persuasion games: (i) whether or not Alice knows Bob’s high-quality product valuation $v$ (complete vs. incomplete information); (ii) whether or not the players know when the game terminates (finite vs. infinite); (iii) whether or not Alice’s messages involve natural language (structured vs. linguistic); (iv) the values of $v$, $p$ and $M$; (v) the value of $T$ in the finite horizon case; and (vi) whether the game is with a long-living Bob or with myopic buyers.
@@
3 Data Collection and Statistical Analysis
In this section, we describe the process of data collection and analysis. We developed a user-friendly game management system to facilitate data collection from the games described in §2. The system is written in Python, designed for ease of use and customization, enabling future researchers to seamlessly collect data. Integrating new language models is straightforward, allowing them to participate in any configurable setup. The interface allows users to effortlessly run data collection across multiple configurations while involving various LLMs in the process. Additionally, the system supports analyzing the collected data to gain insights into the performance and interaction patterns of different models. The system features a simple and intuitive interface, as illustrated in the screenshots provided in Appendix D.1. Prompt samples are described in Appendix D.2.
Table 1: Parameters and their optional values used to define the 1,320 game configurations across bargaining, negotiation, and persuasion game families for data collection. $T=\infty$ means a very large value of $T$, unknown to the players. CI = Complete Information. MA = Textual messages allowed.
@@ 3 Data Collection and Statistical Analysis
In total 384 configurations In total 576 configurations In total 360 configurations
Configurations
@@
Since the game space defined by the parameters presented in §2 is infinite for each of the game families, it is clear that data cannot be collected from all possible games. Therefore, we attempted to cover the game space by selecting diverse values for each of the parameters defining the games, and we collected data from every possible combination generated by these parameters. Table 1 shows the parameters defining the groups from which we collected data. In total, we collected data from 1,320 different configurations: 384 configurations of bargaining, 576 configurations of negotiation, and 360 configurations of persuasion games.
Data
For data collection, we utilized 13 state-of-the-art large LLMs, spanning multiple architectures and vendors. The full list of models is provided in Appendix D.3. The 169 possible pairs (including a language model playing against itself, with attention to the assignment for Alice or Bob) played across 1,320 different configurations. Each of the LLMs played as Alice in 2,500 bargaining games, 2,500 negotiation games, and 1,000 persuasion games, and an equal number of games as Bob. In total, we collected data from 80k games, with full statistics presented in Appendix D.4. To enable comparison with human behavior, we developed an interface for collecting data from games played between LLMs and humans, and gathered data from 3,405 human participants. Details about the human data collection process are provided in Appendix E.
Statistical Analysis
To enable adequate comparisons between models that played in different game configurations, we employ a statistical framework that controls for variation in game structure and player composition. Since each model encountered only a subset of all possible game setups, raw metric values are not directly comparable. To address this, we fit linear regression models that predict each target metric based on the full parameterization of the game and the participating players. This allows us to estimate the independent contribution of each parameter while accounting for confounding effects introduced by the game configuration. In doing so, we obtain normalized estimates of model behavior that can be compared across heterogeneous settings. We train a separate linear regression model for each combination of game family and evaluation metric (introduced in §2). The construction of the feature representation is detailed in Appendix F.1.
The estimated regression coefficients ($\beta$ values) enable us to quantify the impact of each parameter value on the metric relative to a predetermined default value. The list of default values is detailed in Appendix F.2. In addition to the magnitude of the effects, we also report 95% confidence intervals, computed using the standard procedure for linear regression coefficients (see e.g. [49]). Appendix F.3 shows that the linear model performs comparably to state-of-the-art regression models on our prediction tasks, justifying its use in our analyses.
@@
Importantly, metric values presented in this section are calculated by averaging over all possible game configurations in our dataset. These averages are therefore highly sensitive to the particular configurations and the configuration distributions in our dataset. To tailor the benchmark to specific applications, we recommend re-defining the parameter space and their distributions according to the economic context. [^11] For instance, one could ask whether agents’ performance significantly differs in economic environments where inflation is high (translating into lower discount factors, in bargaining games). To evaluate such a scenario, one can simulate games in which the discount factor distribution fits these conditions, and re-evaluate agents’ performance with respect to the new distribution of configurations.
4 Economic and Behavioral Insights
In this section, we present the findings of our analysis, focusing on how different configurations and player compositions affect the outcomes of the games. We examine the general trends observed across all game families and identify specific characteristics that differentiate the performances of language models and human players. We structure our results around the following research questions: Q1: How do the market characteristics, such as information, game horizon, and linguistic communication, affect efficiency and fairness? Q2: How do different LLMs behave in strategic interactions, and which models achieve fair, efficient, and high self-gain outcomes? Q3: How do humans perform compared to LLMs?
Table 2: The effect of the market-defining parameters on the efficiency and the fairness of the game, measured in percentage point improvement relative to the naive parameterization. CI = Complete Information; MA = Textual Messages Allowed; MY=Myopic Buyers; Eff. = Efficiency, Fair. = Fairness; Conf. Int. = Confidence Interval. Bolded values indicate the highest metric values within the confidence interval, while underlined values indicate the lowest metric values.
(a) Bargaining
CI MA T Eff. Fair.
$\infty$ 2.8 4.4
$\infty$ 2.2 2.8
$\infty$ 1.6 0.9
$\infty$
12 -0.2 4.4
12 -0.8 3
12 -1.8 0.9
12 -3.7 -0.8
Conf. Int. $\in$ ±$\pm 0.5$$\pm 0.6$
(b) Negotiation
CI MA T Eff. Fair.
$\infty$ 17.4 0.2
$\infty$ 17.3 0.3
10 16.7 0.1
10 16.3 0.1
10 9.2 -1.2
10 8.3 -2.2
1 6.8 -0.9
1 5.5 -0.3
$\infty$ 5.5 -0.5
$\infty$ 4.1 -1.4
1
1 -2 0.2
Conf. Int. $\in$ ±$\pm 1.7$$\pm 0.3$
(c) Persuasion
CI MA MY Eff. Fair.
(max) (max)
-3.2 -13.4
-5.7 -0.7
-10.9 -6.6
-12.2 -15.6
-13.6 -12.7
-22.4 -22.2
-24.7 -22.7
Conf. Int. $\in$ ±$\pm 2.5$$\pm 2.4$
Table 3: The effect of the Agent on the self gain, for each game family and role in the game.
Family Bargaining Negotiation Persuasion
Model Alice Bob Alice Bob Alice Bob
human 6.6 ± 1.1 -17.1 ± 1 -26.6 ± 1.1 -14.5 ± 1.1 20.5 ± 4.1 54.8 ± 8.9
llama-3.3-70b 1.4 ± 0.7 0.7 ± 0.6 2.5 ± 0.6 4.2 ± 0.6 2.1 ± 2.4 11.2 ± 4.2
claude-3-5-sonnet 1.4 ± 0.7 2.6 ± 0.6 3 ± 0.6 4.5 ± 0.6 3.6 ± 2.4 12.1 ± 4.2
claude-3-7-sonnet 1.4 ± 0.7 2.1 ± 0.6 2.9 ± 0.6 4.4 ± 0.6 4 ± 2.4 11.3 ± 4.2
gpt-4o 1 ± 0.7 0.2 ± 0.6 2.8 ± 0.6 4.2 ± 0.6 4.3 ± 2.4 15.1 ± 4.2
gemini-2.0-flash 0.3 ± 0.7 -0.1 ± 0.6 3.5 ± 0.6 3.7 ± 0.6 -0.1 ± 2.4 21.2 ± 4.2
gemini-1.5-flash 0 ± 0 0 ± 0 0 ± 0 0 ± 0 0 ± 0 0 ± 0
grok-2-1212 -0.3 ± 0.7 0.5 ± 0.6 3 ± 0.7 4.4 ± 0.6 4.6 ± 2.4 11.2 ± 4.2
gpt-4o-mini -0.4 ± 0.7 -1.5 ± 0.6 2.1 ± 0.6 3.5 ± 0.6 0.1 ± 2.4 10.1 ± 4.2
llama-3.1-405b -0.9 ± 0.7 -0.2 ± 0.6 4 ± 0.6 4.7 ± 0.6 -0.1 ± 2.4 21 ± 4.2
mistral-large-2411 -1 ± 0.7 -1.3 ± 0.6 2.5 ± 0.6 5.3 ± 0.6 2.6 ± 2.4 28.3 ± 4.2
gemini-1.5-pro -1 ± 0.7 -1.5 ± 0.6 2.4 ± 0.6 5 ± 0.6 6.3 ± 2.4 25.3 ± 4.2
gemini-2.0-flash-lite -1.4 ± 0.7 -0.3 ± 0.6 2.9 ± 0.6 2.9 ± 0.6 1.1 ± 2.4 9.7 ± 4.2
o3-mini -6.9 ± 0.7 1.4 ± 0.6 2.4 ± 0.6 3.8 ± 0.6 3.4 ± 2.4 24.2 ± 4.2
(a) Alice Gain
1 2 3 4 5 6 7 8 9 A B C D
1. claude-3-5-sonnet
2. claude-3-7-sonnet
3. gemini-1.5-flash
4. gemini-1.5-pro
5. gemini-2.0-flash
6. gemini-2.0-flash-lite
7. gpt-4o
8. gpt-4o-mini
9. grok-2-1212
A. llama-3.1-405b
B. llama-3.3-70b
C. mistral-large-2411
D. o3-mini
0.130.13-0.13- 0.130.020.02-0.02- 0.020.090.090.090.09
(b) Bob’s Gain
1 2 3 4 5 6 7 8 9 A B C D
0.10.1-0.1- 0.10.020.020.020.020.150.150.150.15
Figure 1: The effect of the identities of the two players (rows: Alice, columns: Bob) on self-gain in Bargaining games, reported relative to the mean outcome.
@@ 4 Economic and Behavioral Insights
Table 2 presents the effect of the market-defining parameters on fairness and efficiency, respectively, for each of the game families. From the table, we observe that market conditions have a decisive impact on the fairness and efficiency achieved in the game.
Across all game families, a prolonged interaction between Alice and Bob consistently enhanced both efficiency and fairness. However, the effects of complete information and textual message allowance varied by game type. In bargaining games, complete information reduced both efficiency and fairness, while allowing messages improved both metrics. Conversely, in persuasion games, the pattern was nearly reversed: allowing messages degraded both efficiency and fairness. This held true for scenarios in which Bob was either a casual or a repeat buyer. In contrast, complete information improved fairness in both settings without significantly affecting efficiency. In negotiation games, complete information improved both metrics, whereas message allowance enhanced fairness without impacting efficiency. [^12] We hypothesize that the impact of free language communication differs across game types according to the presence of ground truth. In persuasion games, an objective truth (e.g., product quality) allows agents to build trust through consistent behavior, making free-form language less essential and potentially harmful if it introduces noise. In contrast, bargaining and negotiation games lack an objective ground truth, so language becomes essential for coordination, signaling intentions, and achieving mutually beneficial outcomes.
Interestingly, our results suggest that the effect of complete information on fairness and efficiency may depend on whether message allowance is enabled. In particular, in bargaining games, complete information appears to reduce efficiency and fairness only when linguistic messages are allowed. This interaction is noteworthy because traditional economic models typically abstract away from natural language communication. Our findings suggest that some classic theoretical insights, derived under structured or language-free assumptions, may not be applicable when rich linguistic communication is present. While a comprehensive theoretical investigation is out of scope, we believe this is a promising direction for future work at the intersection of language and economic theory.
LLMs Performance (Q2)
Table 3 presents the agents’ performance as Alice and as Bob in each of the game families in terms of self-gain. Nevertheless, the choice of which LLM to deploy in order to maximize utility is more nuanced than simply selecting the model with the best overall performance. Figure 1 presents the payoffs of Alice (left) and Bob (right) in bargaining games as a function of the identity of the models playing each role. Notably, when Bob is fixed to play with the LLM that yields the highest utility for him across the evaluated LLMs (claude-3.5-sonnet, see Table 3), Alice maximizes her payoff by selecting to play with mistral-large-2411. However, this model generally performs poorly when playing as Alice (ranked 11th out of 13). 131313 One can think of a meta-game in which Alice and Bob (the users) choose LLMs to represent them in an economic game, aiming to maximize their expected utility (over all game parameter realizations). The strategy space consists of available LLMs, and players know only the game family, rather than its specific parameters. Under this setup, the only pure-strategy Nash equilibria are: Bargaining: Alice selects claude-3-5-sonnet and Bob selects claude-3-7-sonnet; Negotiation:Alice selects gemini-2.0-flash and Bob selects gemini-1.5-pro, or Alice selects llama-3.1-405b and Bob selects grok-2-1212; Persuasion: Alice selects gemini-1.5-pro and Bob selects mistral-large-2411.
Figures 21, 22 and 23 in Appendix G.1 show how the pair of models playing negotiation games influenced the efficiency and fairness of the game. These tables suggest that there is no single model that maximizes both efficiency and fairness against all other models, and that compatibility between models plays an important role in shaping these metrics. Furthermore, across all game families, the scenario in which Alice and Bob employed the same LLM neither improved nor degraded the efficiency or fairness of the game, compared to the scenario in which they used different LLMs.
Human Performance (Q3)
Human performance deviates significantly from that of LLMs across the evaluated tasks, in terms of self-gain. In persuasion games, humans consistently outperformed all LLMs by a substantial margin. However, in negotiation games, their self-gain was the lowest among all other LLMs. An intriguing pattern emerged in the bargaining games: while humans performed considerably worse than language models when playing the role of Bob, they significantly outperformed the models when playing as Alice, despite the game’s ostensibly symmetric structure in terms of information. A plausible explanation for this phenomenon lies in a well-documented behavioral bias known as the anchoring effect [45], where the way information is presented can heavily influence human decision-making, even when the substantive content remains unchanged. In our setup, Alice initiates the interaction and anchors the bargaining by making the first offer, effectively setting the frame for the discussion. Unlike LLMs, humans tend to anchor their negotiation behavior to the initial proposal, often in a consistent yet irrational manner. For example, when Bob is a human and Alice is a LLM, the correlation between Alice’s first offer and her final payoff was 0.63, indicating a strong anchoring effect on the human participant. Conversely, when Alice was human and Bob was a LLM, this correlation dropped to just 0.18, suggesting that language models are less influenced by the initial anchor and evaluate offers more rationally than humans, although still not in a fully rational manner.